OpenStudy (anonymous):

A bike has 2 wheels of different sizes. Front wheel is bigger. The smaller wheel rotates 5 times faster than the bigger wheel. The radius of the smaller wheel is 5 inches and the bike moves 3 feet per second. What ia the radius of the bigger wheel? A little help here please. I'm a bit confused on what i am going to do with the '5 times' thingy, i mean, how am i going to epress it and etc :)

OpenStudy (abmon98):

@kmed is your answer 1

OpenStudy (anonymous):

@Abmon u mean, 1in?

OpenStudy (abmon98):

yes

OpenStudy (anonymous):

Yah. I used the formula for circumference. I did it like this: (5)2rpi = 2rpi (5)2rpi = 2(5in)pi r = 1in

OpenStudy (abmon98):

good job @kmeds16 :D

OpenStudy (anonymous):

@Abmon98 thank you! :) so the final answer is really r = 1in right? :)

OpenStudy (abmon98):

your welcome @kmeds16 i am not sure ask more people :)

OpenStudy (anonymous):

@cwrw238 hii, a little help please? :)

OpenStudy (anonymous):

@jim_thompson5910 help me please :)

jimthompson5910 (jim_thompson5910):

What's your question? It looks answered already?

OpenStudy (anonymous):

Looking for more answers & verifications :) most especially from the experts i know :)

OpenStudy (anonymous):

I think my answer is incorrect.

jimthompson5910 (jim_thompson5910):

I'm getting r = 1 as well

jimthompson5910 (jim_thompson5910):

to be more specific, I'm getting 1 ft not 1 inch

jimthompson5910 (jim_thompson5910):

remember the smaller wheel has a radius of 5 inches so the answer has to be larger than 5 inches

OpenStudy (anonymous):

@jim_thompson5910 oh yes yes yes that's what im thinking, so that means it's in feet haha thank you :) so the radius of the bigger wheel is 12 inches coz 1ft is equal to 12 in right

jimthompson5910 (jim_thompson5910):

actually I converted wrong, but I redid it. I'm getting this now v = w1*r1 36 = w1*5 36/5 = w1 w1 = 36/5 w2 = (1/5)*w1 v = w2*r2 v = (1/5)*w1*r2 36 = (1/5)*36/5*r2 (r2)/25 = 1 r2 = 25 So the bigger wheel has a radius of 25 inches

OpenStudy (anonymous):

@jim_thompson5910 ohh i see. Thank you so much! :)

jimthompson5910 (jim_thompson5910):

np